Operation Summary for September 25, 2009 - September 28, 2009 |
The Integrated Luminosity 9/21/09 to 9/27/09 was 27.1 pb-1. Friday Sept. 25 The day shift began with Operations monitoring the Tevatron (TeV) store 7209 (61.27E30) and the Recycler stash (154.7E10), with the Antiproton Source stacking (9.8E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. At 9:43 AM, store 7209 aborted when the TeV RF anode power supply (APS) tripped off due to an interlock fault. There was no quench. RF personnel investigated. Operations prepared CDF and D0 for access at 10:15 AM. Operations reported at 10:47 AM that the TeV RF APS was back online. At 11:14 AM, the Pelletron tripped off with no alarm indication. A Recycler person helped with the investigation. CDF and D0 were out of access by 12:12 PM. Operations put the TeV through a dry squeeze at 12:32 PM. TeV personnel began a study at 1:28 PM. At 3:17 PM, Operations reported that the P1 transfer line was causing stacking problems. MI and Pbar personnel investigated. At 5:04 PM, Operations reported that the Pbar stacking problem was traced to a trim magnet power supply (HTF11) and the Pbar auto tune program. Operations put the TeV into HEP shot setup at 6:10 PM. At 6:47 PM, Operations reported a problem with Booster injection. A knowledgeable Booster person investigated. At 10:01 PM, operators responded to a Linac LCW alarm. They discovered that a LCW hose had blown off a new test source. Operators valved off the water. Operations took the TeV out of shot setup at 10:30 PM. Operations put the TeV through a wet squeeze at 11:13 PM. Saturday Sept. 26 The midnight shift began with Operations monitoring the stash (312.2E10) and stack (47.8E10), and with NuMI and MiniBooNE taking beam. Operations reported at 2:45 AM that Booster tune had increased the beam intensity up to 240E09. Booster tuning continued throughout the rest of the midnight shift. Operations put the TeV into HEP shot setup at 3:33 AM. At 4:02 AM, the TeV suffered a sector F1 quench when a kicker (F17K3) tripped off. The TeV RF APS tripped off due to another interlock fault indication following the quench. Operations called in RF personnel to investigate. At 4:50 AM, Operations reported that the TeV sector F1 cold compressor had tripped off and would not reset. CHL personnel determined that the compressor needed to be replaced. At 7:26 AM, Operations reported that the Pbar autotune program was found turned off; when it was turned back on, it started causing P1 line aborts. Operations called in Pbar personnel to investigate. Operations reported at 9:14 AM that CHL personnel had replaced the F1 cold compressor; cryo system recovery followed, and the pumpdown was complete by 11 AM. Kicker personnel had the F17K3 kicker power supply repaired by 1:30 PM; they replaced a bad control cable. Kicker personnel were sent right back out to look at the problems with a MI-52 kicker (KPS5S). They tried replacing the power supply, but the spare was bad so they put the original supply back in. The P1 line and MI were tuned to compensate for the lower kicker supply voltage until the issue could be addressed Monday morning. Operations reported that the TeV RF personnel had fixed the RF APS. Unfortunately, it tripped again on an interlock fault indication shortly after the RF personnel had left, so Operations called them back in to look into the problem. This happened two more times; they were still working on the problem at the end of the day shift. Booster personnel traced their beam problem to an offset in the sextupole long ramps; once this was fixed, the beam looked good. Operations reported that they stopped transferring antiprotons to the Recycler when the stash reached 454E10. Additional antiprotons were left in the Pbar Accumulator after that. Pbar stacked all shift and had well over 120E10 in the stack by the end of the day shift. The TeV sector B1 house vacuum degraded to the point that it began to affect the cryo system. Operations consulted TeV personnel and then prepared the TeV for access. Vacuum technicians came in to investigate the problem. They traced it to a leaking spool CVI valve in the tunnel. They attached a turbo vacuum pump to the valve and the vacuum improved. Operations turned the TeV on. At 11:50 PM, RF personnel declared victory in their fight against the TeV RF APS, but this was a short-lived victory. The APS tripped off a few minutes later. They resumed their investigation. The Recycler stashed antiprotons when it was appropriate. Sunday Sept. 27 The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. The TeV RF personnel had the RF APS on by 2 AM. Operations had the TeV ramping by 6:52 AM. Operations put the TeV through dry and wet squeezes. Operations checked the orbits and had the TeV in HEP shot setup by 9:45 AM. Operations established store 7219 by 11:30 AM with an average initial luminosity (AIL) of 206.97E30. Pbar personnel began Debuncher cooling studies. Operations terminated store 7219 at 10:31 PM and put the TeV into shot setup. Monday Sept. 28 The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. Operations established store 7220 at 12:20 AM with an AIL of 199.31E30. The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator will allow Pbar to investigate an AP2 PBM problem. In between stores, the controls group will be allowed to change out the TeV sector C3 crate controller. |
|
More Information Comments and Suggestions |
| last modified 09/28/2009 email Fermilab |
| Security, Privacy, Legal | |